Fame is 10 years older than EJ. Five years ago, EJ was two years more than one-fourth as old as Fame. What are their present ages?
Answer by drk(1908)

Let F = fame and E = EJ.
Read each sentence and translate into math.
Fame is 10 years older than EJ.
This translates to:
(i) F = E + 10.
Five years ago, EJ was two years more than one-fourth as old as Fame.
Here you have to reduce both people by 5, -5.
(ii) E - 5 = (1/4)(F - 5) + 2.
We can use substitution of (i) into (ii)
(iii) E - 5 = (1/4)(E + 10 - 5) + 2
Solve for E.
E - 5 = .25E + 2.5 -1.25 + 2
.75E = 8.25
E = 11
F = 21.